OpenStudy (decentnabeel):

A patient is to receive 35 mg of a certain medication at 70 mg per ml of solution.How much of the solution should be given? A=0.35ml B=0.5ml c=2ml D=5ml

OpenStudy (decentnabeel):

@nincompoop @Sachintha

OpenStudy (sachintha):

If a patient needs to take 20mg of medication at 20mg per ml, how many ml should be given?

OpenStudy (sachintha):

Since there are 20mg per ml only 1ml\(\large(\frac{20mg}{20mg.ml^{-1}})\) should be given. Understand?

OpenStudy (sachintha):

Since there is 20mg per ml only 1ml\(\large(\frac{20mg}{20mg.ml^{-1}})\) should be given. Likewise 35mg is needed, but there is 70mg per ml so \(\large(\frac{35mg}{70,g.ml^{-1}})\) should be given.